# The Milky Way in context: The formation of galactic discs and chemical sequences from a cosmological perspective

### v1

@@ 3.1 A diversity of chemical sequences in auriga
We show stellar mass-weighted histograms in the chemical abundance plane for a sample of auriga galaxies in Figure 1, where the full suite is included in an online repository  [^1] We echo the acknowledgment in Grand et al. (2018) that the $\alpha$-abundances in auriga do not necessarily align with the chemistry expected for the MW, and refer the reader to their discussion section where these concerns are explored.. The four chosen examples were selected to exhibit different behaviours; Au-2: with a slanted and narrow track through the chemical abundance plane, reminiscent of the distribution observed in the MW at sub-Solar radii (e.g. figure 6 of Queiroz et al., 2020); Au-14: with a clear bi-modality exhibiting a centrally-dominant high-$\alpha$ sequence, formed in a galaxy that experiences multiple gas-rich mergers on shallow infall trajectories; Au-18: with a clear chemical dichotomy that forms in a realisation with a GSE-like merger occurring at a lookback time of $\tau=9.3\,$Gyr, and a significant metallicity overlap between each chemical sequence; Au-24: with another chemical dichotomy forming in a realisation with a GSE-like merger, this time at $\tau=8.5\,$Gyr. This simulation exhibits a slanted high-$\alpha$ sequence similar to the chemical distribution in the MW, but a less massive low-$\alpha$ sequence that has a narrow metal distribution ($\Delta\rm{[Fe/H]}\approx 0.3$) and hardly any overlap with the metallicity of the high-$\alpha$ sequence, much unlike the MW.
@@
Each panel features a black line, determined by eye, which separates the stars into high-$\alpha$ and low-$\alpha$ selections. Manual selections like these are commonly employed in both observational data and simulated analogues, and our reported results are not sensitive to the exact selection choices. We adjust the selections with each radial bin in order to account for the radial dependence of the chemical sequences in some realisations, which ensures a minimal contamination between the high-$\alpha$ and low-$\alpha$ sequences. The total stellar mass in each selection is indicated in the lower-left corner of each panel. For the purposes of this investigation, we always separate the stars into two selections regardless of whether the chemical distribution adheres to a bi-modal form or not.
@@
We further divide the chemical abundance plane into six radial bins spanning $200<R_{\rm G}/\rm{kpc}<$ 20, which brackets the majority of disc stars at $z=0$. We note that there are also trends with height above and below the disc plane, as seen in the MW (e.g. Nidever et al., 2014; Minchev et al., 2014a; Hayden et al., 2015; Queiroz et al., 2020) and also in auriga at a higher resolution level (Grand et al., 2018). Whilst we do not show these trends explicitly, we find that the proportion of stars in each $\alpha$-sequence is dependent on the $|z|$-height, with stars at higher [Mg/Fe] typically gaining mass dominance at greater $|z|$-heights (especially around and within Solar radii). This is in agreement with observational data of the MW from APOGEE (e.g. Hayden et al., 2015), and is therefore likely to be a general outcome of galaxy formation at MW-mass scales rather than a property that is unique to the MW.
@@
The histograms are coloured by the mean orbital circularity of stars within each pixel (see the definition in Orkney et al. 2023), revealing some interesting features. Firstly, stars in the high-$\alpha$ selection tend to have lower orbital circularities than those in the low-$\alpha$ selection, in-keeping with expectations from our own Galaxy (see for example figure 5 in Chandra et al. 2024 and figure 11 in Khoperskov et al. 2024). Secondly, stars at lower [Fe/H] tend to have lower orbital circularities even at a fixed [Mg/Fe]. This behaviour can also be seen in the MW: see the low-$\alpha$ sequence below a metallicity of approximately $\rm{[Fe/H]}=-0.3$ in figure 6 of Chandra et al. 2024. One interpretation would be that these low-metal stars formed at older times, meaning they have undergone more kinematic perturbations that might heat their orbits, or that they formed at a time when the disc was of a lower orbital circularity. However, both here and in the MW, this wide metallicity spread in the low-$\alpha$ sequence is more accurately a proxy of the underlying radial metallicity gradient than any temporal evolution (i.e. Frankel et al. 2018, 2020, and see also the stellar ages in figure 2 of Patil & Bovy 2024). This is because the stellar disc has kinematically cooled and grown to greater radii during the formation of the low-$\alpha$ sequence, such that the radial metallicity gradient dominates over the temporal metallicity evolution. We will argue in Section 3.3.3 that the lower circularity of these low-metal and low-$\alpha$ stars owes to their formation from recently accreted gases that have not yet fully settled into the rotating gaseous disc.
@@ 3.1.1 Quantifying chemical bi-modalities
From a cursory examination of the full suite, there is clearly a wide variation in the shape and number of $\alpha$-sequences. To aid comparison, we classify the chemical distributions into four groups:
@@
For a more quantitative analysis, we fit a Gaussian profile to the manual selections used in Figure 1, and then estimate the overlap between the two profiles by using the Bhattacharyya distance (Bhattacharyya, 1946) ($D_{\rm B}$):
@@
The figure also demonstrates that radial variation, even between neighbouring radial bins, can have a significant impact on the goodness of the chemical bi-modality. This can be due to both a variation in the mass budget in high-$\alpha$ and low-$\alpha$ sequences, but also due to the changing chemical abundances of stars. An example of the latter effect can be seen in Au-15, where two distinct sequences present at $63<R_{\rm G}/\rm{kpc}<$ 6 begin to converge towards a single mean [Mg/Fe] value at progressively higher radii.
Figure 3: A comparison of four salient properties for the high- and low-$\alpha$ selections described in Section 3.1.1 (for stars in the radial range $200<R_{\rm G}/\rm{kpc}<$ 20). In order; a) The median azimuthal velocity, b) The standard deviation of the velocity in the $z$-direction, c) the summed stellar mass, d) the median of the stellar age. The values for each auriga simulation are provided by unique markers as indicated in the legend. In black, we show expected values for the MW, where kinematic properties are taken from Anguiano et al. (2020), and the high/low-$\alpha$ mass ratio value is taken from Khoperskov et al. (2024).
@@ 3.1.2 Correlation with GSE-like halo debris
Whilst GSE-like mergers do not seem to be a prerequisite in forming multiple chemical sequences, there may still be a requirement for some kind of merger activity. As an example, Au-25 has a particularly quiescent accretion history with the last $>1:30$ mass ratio merger occurring before a lookback time of $\tau=11\,$Gyr. The evolution through the chemical abundance plane is smooth and almost featureless. This is in opposition to results from Prantzos et al. (2023), which finds that a chemical bi-modality can develop from secular evolution even in the absence of mergers. We will return to this idea in Section 3.2.
@@ 3.1.3 Properties of each chemical sequence
We present a comparison of four different properties for the high-$\alpha$ and low-$\alpha$ selections in Figure 3: the median azimuthal velocity, the standard deviation of the velocity in the $z$-direction, the total stellar mass and the median stellar age. These properties are sampled for stars at $z=0$, over the radial range $200<R_{\rm G}/\rm{kpc}<$ 20, and once again exclude stars with halo-like kinematics. The figure shows that, whilst many of the galaxies contain no dichotomous chemical distribution, there are still very predictable overall trends which are shared with the MW. The low-$\alpha$ selections rotate at higher velocities (with Au-26 being the one exception), are kinematically cooler in the vertical direction, and are younger by 2 Gyr or more. These results are an inevitable outcome of the progressive enrichment in MW-mass galaxies, whereby high-$\alpha$ stars are born at higher lookback times. The stellar disc at those times is often rotating at a lower velocity and was kinematically hotter, and/or has been subjected to a greater amount of kinematic heating (e.g. mergers and secular evolution).
@@
We include reference values for the MW in black. The kinematic properties are taken from APOGEE and Gaia DR2 data in Anguiano et al. (2020), who decomposed the Galaxy based on chemical selections of the thick and thin discs. In comparison, the auriga simulations possess stellar discs that are kinematically hotter. This may owe to the numerical limitations inherent to simulations that resolve stellar distributions with discrete mass elements (e.g. particle noise), and cosmological simulations often struggle in reproducing the cooler kinematics observed in the MW (e.g. House et al., 2011; Sanderson et al., 2020). The mass ratio between our high-$\alpha$ and low-$\alpha$ selections has a near random scatter both above and below the $1:1$ dashed-grey line. We compare against the high-$\alpha$ to low-$\alpha$ mass ratio estimate of $4:6$ from the work of Khoperskov et al. (2024), which is based upon an orbit superposition approach. We compare the median age distributions of each sequence with loose predictions that are cited throughout the academic literature. The comparative late formation of the auriga sequences may either be an indication that the MW itself is an outlier with an especially front-loaded assembly history, or that the auriga chemical evolution favours later disc assembly.
@@ 3.2 The time formation of chemical sequences
Figure 4: Each panel shows the star formation histories for each auriga galaxy in units of $\rm{M}_{\odot}\,\rm{yr}^{-1}$. The blue and red histograms are based on the high-$\alpha$ and low-$\alpha$ selections defined in 3.1.1. We stress that not all auriga galaxies display clear chemical bi-modalities, and the type of chemical distribution is indicated by a coloured marker at the top-left of each panel. These selections are subdivided into stacked histograms corresponding to distinct radial bands (based on the stellar positions at $z=0$), as described in the left legend. The total overlapping region is highlighted with a green outline, which shows the contribution of stars that formed coevally. A light-grey histogram shows the additional contribution of stars on halo-like orbits. The pie charts on the right-hand-sides of each panel show the total mass ratios within each radial band. The coeval fraction is once again represented with a green line, but here it is calculated per radial band.
@@
The star formation histories for each auriga galaxy are presented in Figure 4, with two overlaid histograms for the high-$\alpha$ (blue) and low-$\alpha$ (red) selections. These include all stars in the range $200<R_{\rm G}/\rm{kpc}<$ 20 at $z=0$. The lookback times where the two distributions overlap is outlined in green. Whilst many realisations appear to have a significant fraction of coeval stars, these are not necessarily indicative of a coeval formation of the high-$\alpha$ and low-$\alpha$ sequences. In fact, the formation of two chemical sequences in auriga progresses mostly sequentially in all cases. Instead, the fraction of coeval stars are primarily made up of the intermediate-$\alpha$ abundances between the higher and lower sequences. Regardless, it is immediately evident that there are a negligible number of stars in the low-$\alpha$ selections forming at $\tau>9\,$Gyr, similar to what is observed in the MW (though see Borbolato et al. 2025).
@@
These two histograms are subdivided into six stacked histograms, representing the radial regions considered throughout this work. The total mass budgets in each region are shown in the pie charts appended to the right hand side of each panel, which once again includes a green line to show the overlap between high-$\alpha$ and low-$\alpha$ stars (though here, the overlapping fraction is restricted to stars within each radial bin). In most realisations, the stars found in the innermost radial range of $30<R_{\rm G}/\rm{kpc}<$ 3 form either the majority or the plurality of the total stars for both high-$\alpha$ and low-$\alpha$ selections. Yet, it is also common that the overlapping fraction is minimised in this innermost region. This is due to a combination of a few factors. Firstly, the SFR is the most bursty in the inner few kpc, mainly due to its sensitive response to subhalo interactions and inflows. Secondly, gas mixing occurs on shorter timescales, such that there is a narrow chemical abundance range at any one time. Thirdly, there is a lower fraction of stars that migrated from other birth radii where the disc chemistry might be different (which we will show in Section 3.3.5). These effects lead to a higher definition or ‘clumpy’ distribution of stars in the chemical abundance plane.
@@
We stack additional histograms in light-grey, where the thickness represents all remaining stars that do not have disc-like orbits at $z=0$. These can be stars that formed before the development of the stellar disc, stars that naturally formed in regions outside of the disc, or disc stars that have transitioned to halo-like kinematics due to kinematic heating. In some cases, large numbers of disc stars are scattered into the halo during a major merger event. Signs of these major mergers can be seen with sudden and short-lived spikes in the SFR, such as Au-4 at $\tau=2.5\,$Gyrs, Au-19 at $\tau=5\,$Gyrs, or Au-29 at $\tau=4\,$Gyrs. In those most extreme cases, the scattering significantly depletes the mass budget of stars in the high-$\alpha$ sequence.
@@
From inspection of Figure 4, it can be seen that the transitionary epoch between high-$\alpha$ and low-$\alpha$ selections are often preceded by either a period of rising and falling SFR, or a short-lived merger-induced starburst. This is especially evident in realisations with particularly strong dichotomous chemistry (i.e. Au-1, Au-14, Au-18). This behaviour can be explained by considering the interplay between SFR and chemical enrichment: A period of rapid star formation results in a pileup of stars at the current $\alpha$-enrichment of the proto-galaxy. The ensuing $\alpha$-enrichment via rapidly exploding Type-II SNe further encourages the formation of stars with raised $\alpha$ abundances. Both of these mechanisms combined result in the formation of a high-$\alpha$ sequence. Then, as the SFR falls, the iron enrichment from Type-Ia SNe begins to take effect (with a lag of approximately 1 Gyr). The $\alpha$-to-iron ratio drops to a new level, and the low-$\alpha$ sequence begins to form.
@@
We focus on the star formation history of Au-18 in Figure 5, showing only the in-situ stars. As in Figure 4, there is a peak in the star formation history at $\tau=9\,$Gyr that is most prevalent in the innermost radial region, which corresponds to a starburst that is induced by the accretion of a GSE-like merger. This is one example of starburst-induced formation of the chemical thick disc that is described in Grand et al. (2018). However, these stars contribute a much lower mass fraction to the high-$\alpha$ sequence in the Solar neighbourhood ($96<R_{\rm G}/\rm{kpc}<$ 9).
The high-$\alpha$ stars in the Solar neighbourhood arise mostly from a second SFR peak at around $\tau=7\,$Gyr. Curiously, this peak is due to the passage of an extremely minor satellite galaxy with a merger mass ratio of less than $1:100$ at the time of infall, and a peak total mass of just $8\times 10^{9}\,\rm{M}_{\odot}$. It passes beyond the edge of the gas disc at $\tau=7.8\,$Gyr (at which time the disc is mostly confined to within 101010\,10kpc) and briefly disrupts its symmetry, causing star-forming gas to flow beyond the Solar region on one side of the galaxy. This yields a delayed star formation peak that is most prevalent around the Solar neighbourhood, but specifically not at inner radii as with other merger-induced starbursts.
From the decomposition with [Fe/H], there is no obvious sign of a long-lived dilution in the metallicity, as might be expected following the two-infall model of Chiappini et al. (1997, 2001). Instead, there is a sudden and short-lived dilution corresponding to the starburst at $\tau=9\,$Gyr, though this occurs during the formation of the high-$\alpha$ sequence and is not relevant for the later formation of the low-$\alpha$ sequence. Otherwise, there is a gradual progression from metal-poor to metal-rich, with almost a dex of variation at all times and across most radial regions.
The impact of radial migration can be seen when comparing the first and second columns of Figure 5. There is an extinction in the SFR in the two highest radial bins at birth time ($1512<R_{\rm G}/\rm{kpc}<$ 15, $2015<R_{\rm G}/\rm{kpc}<$ 20) between $27>\tau/\rm{Gyr}>$ 2. Before this time period, the compact gas disc is regularly disturbed by merger passages, which act to temporarily propagate star-forming gases to higher radii. For the duration of the extinction time period, the accretion history is extremely quiescent and the gas disc remains undisturbed. The few stars that do form have a low and narrow range of metallicities. By $z=0$, this star formation valley has been repopulated by migrating stars from the inner radial regions. In turn, the range of metallicities broadens and extends to higher values. We will discuss the affects of radial migration further in Section 3.3.5. We note that this star formation valley at the highest radii does not contribute to the formation of the two $\alpha$-sequences in Au-18, as it does in some other auriga simulations (Grand et al., 2018).

### v2

@@ 3.1 A diversity of chemical sequences in auriga
We show stellar mass-weighted histograms in the chemical abundance plane for a sample of auriga galaxies in Figure 1, where the full suite is included in an online repository  [^2] We echo the acknowledgment in grand2018 that the $\alpha$-abundances in auriga do not necessarily align with the chemistry expected for the MW, and refer the reader to their discussion section where these concerns are explored.. The four chosen examples were selected to exhibit different behaviours; Au-2: with a slanted and narrow track through the chemical abundance plane, reminiscent of the distribution observed in the MW at sub-Solar radii (e.g. figure 6 of queiroz2020); Au-14: with a clear bi-modality exhibiting a centrally-dominant high-$\alpha$ sequence, formed in a galaxy that experiences multiple gas-rich mergers on shallow infall trajectories; Au-18: with a clear chemical dichotomy that forms in a realisation with a GSE-like merger occurring at a lookback time of $\tau=9.3\,$Gyr, and a significant metallicity overlap between each chemical sequence; Au-24: with another chemical dichotomy forming in a realisation with a GSE-like merger, this time at $\tau=8.5\,$Gyr. This simulation exhibits a slanted high-$\alpha$ sequence similar to the chemical distribution in the MW, but a less massive low-$\alpha$ sequence that has a narrow metal distribution ($\Delta\rm{[Fe/H]}\approx 0.3$) and hardly any overlap with the metallicity of the high-$\alpha$ sequence, much unlike the MW.
@@
Each panel features a black line, determined by eye, which separates the stars into high-$\alpha$ and low-$\alpha$ selections. Manual selections like these are commonly employed in both observational data and simulated analogues, and our reported results are not sensitive to the exact selection choices. We adjust the selections with each radial bin in order to account for the radial dependence of the chemical sequences in some realisations, which ensures a minimal contamination between the high-$\alpha$ and low-$\alpha$ sequences. The total stellar mass in each selection is indicated in the lower-left corner of each panel. For the purposes of this investigation, we always separate the stars into a high-$\alpha$ and low-$\alpha$ selection even when just one dominant sequence is present. All chemical sequences display some level of granularity in their distribution, and so it is always possible to define an appropriate separation even in the case of single sequences.
@@
We further divide the chemical abundance plane into six radial bins spanning $0<R_{\rm G}/\rm{kpc}<20$, which brackets the majority of disc stars at $z=0$. We note that there are also trends with height above and below the disc plane, as seen in the MW (e.g. nidever2014; minchev2014; hayden2015; queiroz2020) and also in auriga at a higher resolution level (grand2018). Whilst we do not show these trends explicitly, we find that the proportion of stars in each $\alpha$-sequence is dependent on the $|z|$-height, with stars at higher [Mg/Fe] typically gaining mass dominance at greater $|z|$-heights (especially around and within Solar radii). This is in agreement with observational data of the MW from APOGEE (e.g. hayden2015), and is therefore likely to be a general outcome of galaxy formation at MW-mass scales rather than a property that is unique to the MW.
The histograms are coloured by the mean orbital circularity of stars within each pixel (see the definition in orkney2023), revealing some interesting features. Firstly, stars in the high-$\alpha$ selection tend to have lower orbital circularities than those in the low-$\alpha$ selection, in-keeping with expectations from our own Galaxy (see for example figure 5 in chandra2024 and figure 11 in khoperskov2024). This pattern is consistent with a top-down, inside-out formation scenario: the older high-$\alpha$ stars formed when the galaxy was more radially compact but vertically extended, resulting in, on average, less circular orbits. In line with this picture, high-$\alpha$ stars tend to contribute more mass at lower orbital radii than their low-$\alpha$ counterparts. We will show that this is indeed the case for one example simulation in Section 3.3.3.
@@
Secondly, stars at lower [Fe/H] tend to have lower orbital circularities even at a fixed [Mg/Fe]. This behaviour can also be seen in the MW: see the low-$\alpha$ sequence below a metallicity of approximately $\rm{[Fe/H]}=-0.3$ in figure 6 of chandra2024. One interpretation would be that these low-metal stars formed at older times, meaning they have undergone more kinematic perturbations that might heat their orbits, or that they formed at a time when the disc was of a lower orbital circularity. However, both here and in the MW, this wide metallicity spread in the low-$\alpha$ sequence is more accurately a proxy of the underlying radial metallicity gradient than any temporal evolution (i.e. frankel2018; frankel2020; patil2024; cerqui2025). This is because the stellar disc has kinematically cooled and grown to greater radii during the formation of the low-$\alpha$ sequence, such that the radial metallicity gradient dominates over the temporal metallicity evolution. We will argue in Section 3.3.3 that the lower circularity of these low-metal and low-$\alpha$ stars owes to their formation from recently accreted gases that have not yet fully settled into the rotating gaseous disc.
Figure 3: A comparison of four salient properties for the high- and low-$\alpha$ selections described in Section 3.1.1 (for stars in the radial range $0<R_{\rm G}/\rm{kpc}<20$). In order; a) The median azimuthal velocity, b) The standard deviation of the velocity in the $z$-direction, c) the summed stellar mass, d) the median of the stellar age. The values for each auriga simulation are provided by unique markers as indicated in the legend. In black, we show expected values for the MW, where kinematic properties are taken from anguiano2020, and the high/low-$\alpha$ mass ratio value is taken from khoperskov2024.
@@ 3.1.1 Quantifying chemical bi-modalities
From a cursory examination of the full suite, there is clearly a wide variation in the shape and number of $\alpha$-sequences. To aid comparison, we manually classify the chemical distributions into four groups:
@@
For a more quantitative analysis, we fit a Gaussian profile to the manual selections used in Figure 1, and then estimate the overlap between the two profiles by using the Bhattacharyya distance (Bhattacharyya1946) ($D_{\rm B}$):
@@
We use the assumption that the chemical sequences are well-described by Gaussian profiles, despite the lack of direct physical justification for this choice. Nevertheless, the Gaussian models provide good fits in nearly all cases (e.g., see the histograms along the right $y$-axis in Figure 2). We also remind the reader that our manual selections divide the chemical plane into the most representative high-$\alpha$ and low-$\alpha$ populations, even in cases where the distribution appears predominantly uni-modal or smeared. In the case of smeared distributions, the Gaussian fits are far poorer.
@@
The figure also demonstrates that radial variation, even between neighbouring radial bins, can have a significant impact on the goodness of the chemical bi-modality. This can be due to both a variation in the mass budget in high-$\alpha$ and low-$\alpha$ sequences, but also due to the changing chemical abundances of stars. An example of the latter effect can be seen in Au-15, where two distinct sequences present at $3<R_{\rm G}/\rm{kpc}<6$ begin to converge towards a single mean [Mg/Fe] value at progressively higher radii.
@@ 3.1.2 Correlation with GSE-like halo debris
Whilst GSE-like mergers do not seem to be a prerequisite in forming multiple chemical sequences, there may still be a requirement for some kind of merger activity. As an example, Au-25 has a particularly quiescent accretion history with the last >1:30>1:30 mass ratio merger occurring before a lookback time of $\tau=11\,$Gyr. The evolution through the chemical abundance plane is smooth and almost featureless. This is in opposition to results from prantzos2023, which finds that a chemical bi-modality can develop from secular evolution even in the absence of mergers. We will return to this idea in Section 3.2.
@@ 3.1.3 Properties of each chemical sequence
Figure 4: Included are a selection of auriga realisations with bi-modal (column 1), uni-modal (column 2) and smeared chemistry (column 3). The entire suite can be seen in Appendix A. Each panel shows the star formation history in units of $\rm{M}_{\odot}\,\rm{yr}^{-1}$. The blue and red histograms are based on the high-$\alpha$ and low-$\alpha$ selections defined in 3.1.1. These selections are subdivided into stacked histograms corresponding to distinct radial bands (based on the stellar positions at $z=0$), as described in the left legend. The total overlapping region is highlighted with a green outline, which shows the contribution of stars that formed coevally. A light-grey histogram shows the additional contribution of stars on halo-like orbits. The pie charts on the right-hand-sides of each panel show the total mass ratios within each radial band. The coeval fraction is once again represented with a green line, but here it is calculated per radial band.
@@
We present a comparison of four different properties for the high-$\alpha$ and low-$\alpha$ selections in Figure 3: a) the median azimuthal velocity, b) the standard deviation of the velocity in the $z$-direction, c) the total stellar mass and d) the median stellar age. These properties are sampled for stars at $z=0$, over the radial range $0<R_{\rm G}/\rm{kpc}<20$, and once again exclude stars with halo-like kinematics following the methods described in 2. The figure shows that, whilst many of the galaxies contain no dichotomous chemical distribution, there are still very predictable overall trends which are shared with the MW. The low-$\alpha$ selections rotate at higher velocities (with Au-26 being the one exception), are kinematically cooler in the vertical direction, and are younger by 2 Gyr or more. These results are an inevitable outcome of the progressive enrichment in MW-mass galaxies, whereby high-$\alpha$ stars are born at higher lookback times. The stellar disc at those times is often rotating at a lower velocity and was kinematically hotter, and/or has been subjected to a greater amount of kinematic heating (e.g. mergers and secular evolution).
@@
We include reference values for the MW in black. The kinematic properties are taken from APOGEE and Gaia DR2 data in anguiano2020, who decomposed the Galaxy based on chemical selections of the thick and thin discs. In comparison, the auriga simulations possess stellar discs that are kinematically hotter. This may owe to the numerical limitations inherent to simulations that resolve stellar distributions with discrete mass elements (e.g. particle noise), and cosmological simulations often struggle in reproducing the cooler kinematics observed in the MW (e.g. house2011; sanderson2020). The mass ratio between our high-$\alpha$ and low-$\alpha$ selections has a near random scatter both above and below the $1:1$ dashed-grey line. We compare against the high-$\alpha$ to low-$\alpha$ mass ratio estimate of $4:6$ from the work of khoperskov2024, which is based upon an orbit superposition approach. We compare the median age distributions of each sequence with loose predictions that are cited throughout the academic literature. The comparative late formation of the auriga sequences may either be an indication that the MW itself is an outlier with an especially front-loaded assembly history, or that the auriga chemical evolution favours later disc assembly.
A key result in panel d) is the wide variation of formation times for both high-$\alpha$ and low-$\alpha$ sequences. The formation of a bi-modality can then be considered to be the time when a low-$\alpha$ sequence has grown a significant mass fraction, which is qualitatively similar to the median low-$\alpha$ formation times shown here. This highlights that chemical bi-modalities depend on the unique assembly and star formation history of each galaxy rather than intrinsic universal formation epochs, a topic we explore in the next section.
@@ 3.2 The time formation of chemical sequences
The star formation histories for a selection of auriga galaxies are presented in Figure 4, where we consider examples that have bi-modal, uni-modal and smeared chemistry. The full suite can be found in Appendix A. Each star formation history is represented with two overlaid histograms for the high-$\alpha$ (blue) and low-$\alpha$ (red) selections, including all stars in the range $0<R_{\rm G}/\rm{kpc}<20$ at $z=0$.
@@
The lookback times where the two distributions overlap is outlined in green. Whilst many realisations appear to have a significant fraction of coeval stars, these are not necessarily indicative of a coeval formation of the high-$\alpha$ and low-$\alpha$ sequences. In fact, the formation of two chemical sequences in auriga progresses mostly sequentially in all cases. Instead, the fraction of coeval stars are primarily made up of the intermediate-$\alpha$ abundances between the higher and lower sequences. Regardless, it is immediately evident that there are a negligible number of stars in the low-$\alpha$ selections forming at $\tau>9\,$Gyr, similar to what is observed in the MW (though see borbolato2025).
@@
These two histograms are subdivided into six stacked histograms, representing the radial regions considered throughout this work. The total mass budgets in each region are shown in the pie charts appended to the right hand side of each panel, which once again includes a green line to show the overlap between high-$\alpha$ and low-$\alpha$ stars (though here, the overlapping fraction is restricted to stars within each radial bin). In most cases, the stars found in the innermost radial range of $0<R_{\rm G}/\rm{kpc}<3$ form either the majority or the plurality of the total stars for both high-$\alpha$ and low-$\alpha$ selections (though with the low-$\alpha$ stars in ‘smeared chemistry’ examples proving an exception). Yet, it is also common that the overlapping fraction is minimised in this innermost region. This is due to a combination of a few factors. Firstly, the SFR is the most bursty in the inner few kpc, mainly due to its sensitive response to subhalo interactions and inflows. Secondly, gas mixing in inner regions occurs on shorter timescales, such that there is a narrow chemical abundance range at any one time. Thirdly, there is a lower fraction of stars that migrated from other birth radii where the disc chemistry might be different (which we will show in Section 3.3.5). These effects lead to a higher-definition or ‘clumpy’ distribution of stars in the chemical abundance plane.
@@
We stack additional histograms in light-grey, where the thickness represents all remaining stars that do not have disc-like orbits at $z=0$. These can be stars that formed before the development of the stellar disc, stars that naturally formed in regions outside of the disc, or disc stars that have transitioned to halo-like kinematics due to kinematic heating. In some cases, large numbers of disc stars are scattered into the halo and onto higher radial orbits during a major merger event. Signs of these major mergers can be seen with sudden and short-lived spikes in the SFR, such as Au-27 at $\tau=9.5\,$Gyrs. In the most extreme cases, the scattering depletes the mass budget of stars in the high-$\alpha$ sequence, which in turn reduces the strength of any potential bi-modality.
@@
The transitionary epoch between high-$\alpha$ and low-$\alpha$ selections are often preceded by either a period of rising and falling SFR, or a short-lived merger-induced starburst. This is especially evident in realisations with strong dichotomous chemistry, i.e. the ‘bi-modal’ examples. We now explain how the interplay between SFR and chemical enrichment can contribute to the formation of a bi-modality: A period of rapid star formation results in a pileup of stars at the current $\alpha$-enrichment of the proto-galaxy. The ensuing $\alpha$-enrichment via rapidly exploding Type-II SNe further encourages the formation of stars with raised $\alpha$/Fe] abundances. Both of these mechanisms combined result in the formation of a high-$\alpha$ sequence. Then, as the SFR falls, the iron enrichment from Type-Ia SNe begins to take effect (with a lag of approximately 1 Gyr). The $\alpha$-to-iron ratio drops to a new level, and the low-$\alpha$ sequence begins to form. This is, in essence, a description of the $\alpha$-knee as in tinsley1979; tinsley1980; mcwilliam1997.
Whilst some ‘uni-modal’ realisations share this SFR feature, there are circumstances that have inhibited the formation of a bi-modality. For example, in Au-27 the merger-induced starburst at $\tau=9.5\,$Gyr occurred when the host galaxy had not grown much beyond 3 kpc. As a result, some chemical dichotomy is present in the innermost regions but does not reach as far as the Solar radius. The same scenario can be seen occurring for Au-12 in Appendix A. There is no significant rise and fall in the SFR outside of the galactic centre, and this is shown to greater depth in Section 3.2.1.
Across the entire suite, there is a great variation in the star formation history for realisations with bi-modal and uni-modal chemistry. Certainly, there are multiple routes to forming different chemical morphologies. In contrast, the ‘smeared chemistry’ examples display very similar characteristic histories: an early radial growth of the disc, and then a gradual reduction in the SFR until the present day. This early radial growth is a consequence of high angular momenta accretions from the cosmic web during the early galactic assembly. The relatively steady SFR over all radial bins, and the scarcity of disruptive radial merger events when compared with the full suite, results in a smooth chemical enrichment through time. As such, these realisations lack a chemical bi-modality.
@@
We focus on the star formation history of Au-18 in Figure 5, showing only the in-situ stars. As in Figure 4, there is a peak in the star formation history at $\tau=9\,$Gyr that is most prevalent in the innermost radial region, which corresponds to a starburst that is induced by the accretion of a GSE-like merger. This is one example of starburst-induced formation of the chemical thick disc that is described in grand2018. However, these stars contribute a much lower mass fraction to the high-$\alpha$ sequence in the Solar neighbourhood ($6<R_{\rm G}/\rm{kpc}<9$).
The high-$\alpha$ stars in the Solar neighbourhood arise mostly from a second SFR peak at around $\tau=7\,$Gyr. Curiously, this peak is due to the passage of an extremely minor satellite galaxy with a merger mass ratio of less than $1:100$ at the time of infall, and a peak total mass of just $8\times 10^{9}\,\rm{M}_{\odot}$. It passes beyond the edge of the gas disc at $\tau=7.8\,$Gyr (at which time the disc is mostly confined to within 1010\,kpc) and briefly disrupts its symmetry, causing star-forming gas to flow beyond the Solar region on one side of the galaxy. This yields a delayed star formation peak that is most prevalent around the Solar neighbourhood, but specifically not at inner radii as with other merger-induced starbursts.
From the decomposition with [Fe/H], there is no obvious sign of a long-lived dilution in the metallicity, as might be expected following the two-infall model of chiappini1997; chiappini2001. Instead, there is a sudden and short-lived dilution corresponding to the starburst at $\tau=9\,$Gyr, though this occurs during the formation of the high-$\alpha$ sequence and is not relevant for the later formation of the low-$\alpha$ sequence. Otherwise, there is a gradual progression from metal-poor to metal-rich, with almost a dex of variation at all times and across most radial regions.
The impact of radial migration can be seen when comparing the first and second columns of Figure 5. There is an extinction in the SFR in the two highest radial bins at birth time ($12<R_{\rm G}/\rm{kpc}<15$, $15<R_{\rm G}/\rm{kpc}<20$) between $7>\tau/\rm{Gyr}>2$. Before this time period, the compact gas disc is regularly disturbed by merger passages, which act to temporarily propagate star-forming gases to higher radii. For the duration of the extinction time period, the accretion history is extremely quiescent and the gas disc remains undisturbed. The few stars that do form have a low and narrow range of metallicities. By $z=0$, this star formation valley has been repopulated by migrating stars from the inner radial regions. In turn, the range of metallicities broadens and extends to higher values. We will discuss the affects of radial migration further in Section 3.3.5. We note that this star formation valley at the highest radii does not contribute to the formation of the two $\alpha$-sequences in Au-18, as it does in some other auriga simulations (grand2018).